verify the MLE of common p under Ho

mle: \hat{p}=\frac{X+Y}{m+n}
Answer #1

11 x +4 - о log Lo Әр mtn-x-y 1-р ( 12 )(t-p) - (m+n-x-3) = - та - тръпрхK + х = 0 - U (х+3) - (m+n) р-о h+x=& (utul R+x=d mtany query, please ask